# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v10

@@ 3.4 Data Description
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^44] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995). As this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^45] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts rarely occurred among the THBS households. Although family size was included in all regressions, to be conservative, it had little effect on the results. [^46] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and were absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
@@ Results
(1) Semi-monthly Panels (2) Adj. Monthly Panels (3) Unadj. Monthly Panels
Disposable income Obs. Disposable income Obs. Disposable income Obs.
Coef. Std. error Coef. Std. error Coef. Std. error
Total consumption 0.358 [0.033]*** 278 0.433 [0.059]*** 123 0.261 [0.054]*** 141
      Food 0.136 [0.030]*** 259 0.293 [0.066]*** 114 0.100 [0.072] 132
      Housing 0.099 [0.415] 134 0.013 [0.084] 105 -0.263 [0.167] 124
      Utilities -0.003 [0.138] 223 0.343 [0.236] 110 0.197 [0.334] 130
      Furniture 0.426 [0.223]* 169 -0.114 [0.706] 98 0.250 [0.339] 112
      Clothes 0.378 [0.190]* 238 0.347 [0.424] 112 0.806 [0.262]*** 132
      Education 0.101 [0.051]* 236 0.162 [0.128] 110 0.069 [0.146] 128
      Medical expenses 0.039 [0.066] 258 0.194 [0.120] 114 -0.128 [0.094] 132
      Entertainment expenses 0.307 [0.084]*** 250 0.716 [0.181]*** 113 0.194 [0.153] 132
      Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97 0.362 [0.291] 115
      Gifts 0.523 [0.106]*** 234 0.469 [0.197]** 113 0.622 [0.306]* 130
      Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110 0.246 [0.267] 126
@@
Notes: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories. I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable. [^59] For example, the estimate for housing was negative. The households paid a fixed rent amount monthly. Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these months may have created a strong but false positive correlation. This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
@@
4.3 Event-Study Analysis: Response to Adverse Health Shocks
@@ 5 Credit System in Local Economy
The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Therefore, systematic information on credit purchases is unavailable. Fortunately, however, I found that some official survey reports documented information on this credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^71] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^72] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^73] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^74] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^75] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^77] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
Table 10: Volatility of Earnings and Use of Credit Purchases
Dependent Variable: Indicator Variable for
Non-CP Users Non-CP + Low-CP Users
(1) (2) (3) (4) (5) (6)
CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
Model LPM LPM Probit LPM LPM Probit
Number of THBS households 33 33 33 33 33 33
Family Size Yes Yes Yes Yes Yes Yes
Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
Head’s age No Yes Yes No Yes Yes
Head’s earnings No Yes Yes No Yes Yes
Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
Column (1) of Table 10 presents the result for Equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases. [^78] Despite this, low-income households might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
The results so far indicate that retailers were generous in considering the earnings volatility risk. This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^79] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^80] The repayment period in credit purchases among retailers and wholesalers (footnote 74) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
@@ 6 Conclusion
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks.
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916898989This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.909090The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^91] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.

### v11

@@ 3.4 Data Description
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^43] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995). As this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^44] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts rarely occurred among the THBS households. Although family size was included in all regressions, to be conservative, it had little effect on the results. [^45] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. I then used the share of children below primary school age range as an alternative time-varying preference shock variable. Including the child share control variable does not influence the results, and the estimates are materially similar to those from my baseline specification (Online Appendix Table B.6). Panel C of Table 3 shows the summary statistics on the family size variable.
@@ Results
(1) Semi-monthly Panels (2) Adjusted Monthly Panels
Disposable income Obs. Disposable income Obs.
Coef. Std. error Coef. Std. error
Total consumption 0.358 [0.033]*** 278 0.433 [0.059]*** 123
      Food 0.136 [0.030]*** 259 0.293 [0.066]*** 114
      Housing 0.099 [0.415] 134 0.013 [0.084] 105
      Utilities -0.003 [0.138] 223 0.343 [0.236] 110
      Furniture 0.426 [0.223]* 169 -0.114 [0.706] 98
      Clothes 0.378 [0.190]* 238 0.347 [0.424] 112
      Education 0.101 [0.051]* 236 0.162 [0.128] 110
      Medical expenses 0.039 [0.066] 258 0.194 [0.120] 114
      Entertainment expenses 0.307 [0.084]*** 250 0.716 [0.181]*** 113
      Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97
      Gifts 0.523 [0.106]*** 234 0.469 [0.197]** 113
      Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110
@@
Notes: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed. Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
4.3 Response to Adverse Health Shocks
@@ 5 Credit System in Local Economy
The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Fortunately, however, I found that some official survey reports documented information on this credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^69] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). Thus, credit purchases in Tokyo were neither region- nor religion-specific customs. A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^70] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^71] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^72] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^73] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^75] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large. Online Appendix C.5 summarizes the results. Shortly, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant in different specifications. This result remains unchanged when alternative thresholds for the dependent indicator variable are used to include the households with less than the first quartile of the average credit purchases. This suggests that retailers may be generous in considering the volatility of earnings when providing credit to factory workers, which is consistent with the retailers’ incentive to secure their local customers.
@@
This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^76] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^77] The repayment period in credit purchases among retailers and wholesalers (footnote 72) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
@@ 6 Conclusion
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks. A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916868686This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.878787The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^88] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.